A population was scored for the MN blood type locus. There were 97 Type M, 447 Type MN, and 195 Type N individuals in the population. Calculate the frequency of the M allele in this population.Round your answer properly to four decimal digits.
Answer #1
  • Total population = 97+ 447+195= 739

  • Since, each individual will have two alleles (MM,MN or NN) determining the blood group-

Total number of alleles in the population = 2 x 739 =1478 = Total or T

  • Individuals with M type or MM genotype = 97 =D (D for dominant)

  • Individuals with M type or MN genotype = 447 =H (H for Heterozygous)

  • Individuals with M type or NN genotype = 195 = R (R for recessive)

Since, M allele is present in both genotype MM and MN, the allele frequency =

  • f(M) = (2D + H)/ T
  • f(M) = {(2 x 97) + 447}/ 1478 = 0.4337 (Aprox)
  • Allele frequency for M allele in the population = 0.4337